# Recent Advances in Discrete Speech Tokens: A Review

### v2

@@ II-B3 Finite Scalar Quantization
The quantized values are thus integers ranging from $-\lfloor L/2\rfloor$ to $\lfloor L/2\rfloor$ [^2] Following [58], this is the symmetric case for $L$ being odd. When $L$ is even, there is an offset before rounding to obtain asymmetric quantized values.. For a $d$-dimensional vector $\bm{x}$, there are $L^{d}$ possible quantization outcomes. Hence, FSQ usually requires a much smaller hidden dimension than VQ. STE is applied to pass gradients. As quantization is simply done via rounding to integers, there is no explicit codebooks associated with the FSQ process. FSQ is reported to have better codebook usage [^3] Although there is no longer a codebook associated with code-vectors, codebook usage can still be measured among all possible $V=L^{d}$ outcomes. especially for a large $V$ compared to VQ methods, without the need for auxiliary losses.
@@ II-B4 Other VQ Tricks
Figure 3: Diagram of GVQ (left) and RVQ (right).
@@
    Grouped VQ (GVQ), also known as product quantization [59]. It groups the input vector $\bm{x}$ by dimensions and apply VQ on different parts of $\bm{x}$ independently. They can have different or shared codebooks. The VQ outputs are then concatenated along dimensions to match that of $\bm{x}$. For instance, GVQ is used in neural word embeddings [60] and speech self-supervised learning models [31, 32] to achieve efficient quantization.
@@
    Residual VQ (RVQ), also known as multi-stage quantization [61]. It adopts a serial approach that iteratively quantizes the residual of the last quantizer. Similar to GVQ, RVQ also has multiple quantizers. For the $i$-th quantizer $q_{i}$ with input $\bm{x}_{i}$ and output code-vector $\bm{c}_{k}$, the residual is defined as $\bm{x}_{i+1}=\bm{x}_{i}-\bm{c}_{k}$. The outputs from all $q_{i}$ are finally summed as the quantized result of $\bm{x}$. In this way, information in the codebooks is supposed to follow a coarse-to-fine order, and more details in the original $\bm{x}$ can be preserved than a plain quantizer. It is used in various speech codecs [62, 63, 19], for instance.
GVQ and RVQ can be flexibly combined to form GRVQ [64] that applies RVQ on each GVQ branch for better codebook utilization. A network can also contain multiple VQ modules at different places, like cross-scale VQ (CSVQ) [65] where every decoder layer has a quantizer inside.
Note that RVQ naturally produces an order of importance in residual layers, while all quantizers in GVQ are equally important. Such order of importance can also be enforced in GVQ by a “nested dropout” trick [66].
@@ III Speech Tokenization Methods: Acoustic Tokens
TABLE I: A summary of famous acoustic speech tokens (neural speech codecs). Italic “C,T,U” denote CNN, Transformer or U-Net-based generator architecture in Fig.5. Symbols ‘/’ and ‘-’ denote “or” and “to” for different model versions, and “+” means different configurations in different VQ streams in a single model. $Q,F,V$ mean number of quantizers, frame rate and vocabulary size of each quantizer respectively. For example, “$Q=2$, $V$=8192+($2^{12}$-$2^{15}$)” in SemantiCodec means one of the two VQ streams has 8192 possible codes, and the other can vary from $2^{12}$ to $2^{15}$ in different configurations. Bitrates are computed by $\frac{1}{1000}\sum_{i=1}^{Q}F_{i}\lceil\log_{2}V_{i}\rceil$ kbps, without entropy coding.
General-purpose acoustic tokens
SoundStream [62] VQ-GAN (C) 24 RVQ max 24 75 1024 max 18.00
EnCodec [63] VQ-GAN (C) 24 RVQ max 32 75 1024 max 24.00
TF-Codec [57] VQ-GAN (C) 16 GVQ 3-32 25 512 / 1024 0.68-8.00
Disen-TF-Codec [21] VQ-GAN (C) 16 GVQ 2 / 6 25 256 / 1024 0.40 / 1.50
AudioDec [67] VQ-VAE (C)+GAN 48 RVQ 8 160 1024 12.80
HiFi-Codec[64] VQ-GAN (C) 16 / 24 GRVQ 4 50-100 1024 2.00-4.00
DAC [19] VQ-GAN (C) 44.1 RVQ 9 86 1024 7.74
LaDiffCodec [17] Latent diffusion 16 RVQ 3 / 6 50 1024 1.50 / 3.00
FreqCodec [68] VQ-GAN (C) 16 RVQ max 32 50 1024 max 16.00
TiCodec [22] VQ-GAN (C) 24 RVQ, GVQ 1-4 75 1024 0.75-3.00
APCodec [69] VQ-GAN (C) 48 RVQ 4 150 1024 6.00
SRCodec [70] VQ-GAN (C) 16 GRVQ 2-8 50 512+1024 0.95-3.80
SQ-Codec [71] VQ-GAN (C) 16 FSQ 32 50 19 8.00
Single-Codec [72] VQ-GAN (T+C) 24 VQ 1 23 8192 0.30
ESC [73] VQ-GAN (U) 16 GVQ max 18 50 1024 max 9.00
CoFi-Codec [74] VQ-GAN (U) 16 GVQ 3 8.33+25+50 16384 1.17
HILCodec [75] VQ-GAN (C) 24 RVQ 2-12 75 1024 1.50-9.00
SuperCodec [76] VQ-GAN (C) 16 RVQ 2-12 50 1024 1.00-6.00
SNAC [23] VQ-GAN (C) 24 RVQ 3 12+23+47 4096 0.98
WavTokenizer [77] VQ-GAN (C) 24 VQ 1 40 / 75 4096 0.48 / 0.90
BigCodec [78] VQ-GAN (C) 16 VQ 1 80 8192 1.04
LFSC [79] VQ-GAN (C) 22.05 FSQ 8 21.5 2016 1.89
NDVQ [80] VQ-GAN (C) 24 RVQ max 32 75 1024 max 24.00
VRVQ [24] VQ-GAN (C) 44.1 RVQ 8 86 1024 0.26 + max 6.89
TS3-Codec [20] VQ-GAN (T) 16 VQ 1 40 / 50 $2^{16}$ / $2^{17}$ 0.64-0.85
Stable-Codec [81] VQ-GAN (T) 16 FSQ 6 / 12 25 5 / 6 0.40 / 0.70
FreeCodec [82] VQ-GAN (C+T) 16 VQ 1+1 50+7 256 0.45
Mixed-objective acoustic tokens: semantic distillation
Siahkoohi et al. [83] VQ-GAN (C) 16 RVQ 2+1 / 2+2 / 6 25+50 64 0.60 / 0.90 / 1.80
SpeechTokenizer [9] VQ-GAN (C) 16 RVQ 8 50 1024 4.00
SemantiCodec [18] Latent diffusion 16 VQ 2 12.5-50 8192+($2^{12}$-$2^{15}$) 0.31-1.40
LLM-Codec [25] VQ-GAN (C) 16 RVQ 3 8.33+16.67+33.33 3248+32000+32000 0.85
X-Codec [26] VQ-GAN (C) 16 RVQ max 8 50 1024 max 4.00
SoCodec [84] VQ-GAN (C) 16 GVQ 1 / 4 / 8 25 / 8.3 / 4.2 16384 0.35 / 0.47
Mimi [10] VQ-GAN (C+T) 24 RVQ 8 12.5 2048 1.10
X-Codec 2.0 [85] VQ-GAN (C+T) 16 FSQ 8 50 4 0.80
Mixed-objective acoustic tokens: disentanglement
SSVC [27] VQ-GAN (C) 24 RVQ 4 50 512 1.80
PromptCodec [86] VQ-GAN (C) 24 GRVQ 1-4 75 1024 0.75-3.00
FACodec [28] VQ-GAN (C) 16 RVQ 1+2+3 80 1024 4.80
LSCodec [29] VQ-VAE (C+T)+GAN 24 VQ 1 25 / 50 1024 / 300 0.25 / 0.45
SD-Codec [30] VQ-GAN (C) 16 RVQ 12 50 1024 6.00
Acoustic tokens, also known as speech codecs, refer to the discrete representations optimized mainly for signal compression and reconstruction. The audio codec technology arises long ago. Traditional codecs, including MP3 [87], Opus [88] and EVS [89], typically take advantage of signal processing algorithms to improve quality and lower the bitrate.
In the deep learning era, numerous codec models based on neural networks have emerged. These models typically consist of an encoder that compresses speech signals and a decoder that reconstructs the speech signals, with a quantizer situated between the two. The quantizer is also parameterized and jointly trained with the whole network in an end-to-end manner. The codebook indices produced by the quantizer are referred to as acoustic tokens. To improve the representation ability of discrete VQ spaces and thus obtain better codec performance, RVQ, GVQ, GRVQ and FSQ tricks are commonly applied in the quantization module.
We list the VQ method, number of quantizers $Q$, frame rate $F$, vocabulary size $V$ for each quantizer, and the resulting bitrate of existing neural acoustic speech tokens in Table.I.
@@ III-A Model Architectures
Although acoustic codec models differ from one to one regarding their purposes, most of them share a similar encoder-quantizer-decoder framework. With audio clip $\bm{x}$ that can either be time-domain sampling points, frequency-domain features or even other machine learning features, an encoder $f_{\theta}(\cdot)$ transforms it to $f_{\theta}(\bm{x})$ in a continuous latent vector space. The encoder $f_{\theta}(\cdot)$ will usually perform downsampling to reduce the temporal length of the input signals, especially for waveform inputs. A VQ module $q_{\phi}(\cdot)$ discretizes $f_{\theta}(\bm{x})$ into tokens and corresponding codebook vectors $\bm{c}$. A decoder $g_{\psi}(\cdot)$ then uses $\bm{c}$ to reconstruct $\hat{\bm{x}}$, and a certain distance metric of $d(\bm{x},\hat{\bm{x}})$ is usually optimized. There are two major paradigms for designing the encoder, decoder, and quantizers, which can be summarized as diagrams in Fig.4.
Figure 5: Major generator (VQ-VAE) architectures of VQ-GAN-based acoustic tokens. “Q.” and “Trans” are short for quantizer and Transformer, respectively.
@@ III-A1 VQ-GAN
VQ-GAN [90] is a very commonly adopted framework of acoustic tokens that trains a VQ-VAE with GAN objectives. Besides the original reconstruction and VQ objectives in a VQ-VAE, VQ-GAN uses discriminators $d_{\xi}(\bm{x},\hat{\bm{x}})$ to distinguish real and reconstructed data that adversarially train the generator network composed of $f_{\theta},q_{\phi}$, and $g_{\psi}$. In acoustic tokens, there are usually multiple discriminators, e.g. multi-resolution and multi-scale STFT discriminators from the neural vocoder researches [91, 92]. The generator architecture of VQ-GAN-based acoustic tokens has multiple choices, with the three most representative ones visualized in Fig.5: CNN-based, Transformer-based, and U-Net-based.
@@
The CNN-based generator is the most widely used architecture so far in acoustic tokens. SoundStream [62] and EnCodec [63] are two famous early neural acoustic tokens that operate in an end-to-end VQ-GAN manner. They receive time-domain waveforms as inputs and directly reconstruct waveforms. Their encoder and decoder have a mirrored architecture to perform down and up-samplings. In SoundStream, the encoder and decoder are purely constructed by convolutional neural networks (CNNs) while EnCodec augments them with an LSTM. The CNN encoder down-samples the waveform to a high-dimensional embedding sequence, whose frame rate is determined by the sampling rate, CNN kernel sizes and strides at a fixed ratio. The continuous embeddings are passed to an RVQ quantizer, and the quantized vectors are summed before being transformed to the waveform domain by the CNN decoder. The training criteria include reconstruction loss (in the time and frequency domain), adversarial loss, feature matching loss, and quantization losses for RVQ layers. To allow for a flexible choice of bitrates, structured dropout is adopted where the number of codebooks in the RVQ module can be randomly chosen [62], such that only a portion of quantizers in front are activated during training. The acoustic tokens can consequently reside in variable bitrates depending on the chosen number of RVQ quantizers. The inputs and outputs of the codec model can also be frequency-domain features like magnitude and phase spectra for reducing computation burden [68]. There, the convolution kernels are typically 2D instead of 1D in the time-domain codecs.
@@
Later, Transformers [3] have been adopted, e.g. in Single-Codec [72] and Mimi [10]. They can be directly applied to frequency-domain inputs and outputs. When operating on waveform-domain inputs or outputs, a CNN [10] or patchifying [20, 81] operation is usually added before or after the Transformer blocks. In Mimi, a shallow Transformer layer is added after the CNN-based encoder, and vice versa in its decoder. Recently, some propose to use purely Transformer-based backbone and discard the CNN blocks, e.g. TS3-Codec [20]. As Transformers demonstrate superior modeling ability and scaling property, these works prove to outperform CNN-based codecs either with less computation [20] or larger scale [81]. However, to ensure stream-ability, an attention mask should be employed [10]. The encoder and decoder can also be designed to be different. For example, Single-Codec [72] uses Conformer [93] encoder and CNN decoder, while LSCodec [29] uses the reverse configuration.
Though RVQ or GVQ is usually applied, most acoustic tokens contain only one quantization module as a whole. However, there are also U-Net-based codecs where multiple quantizers are employed, e.g. CoFi-Codec [74] and ESC [73]. Each sub-encoder or decoder in the U-Net can be a CNN or Transformer. This offers a more flexible control of the resolution of each VQ stream (Section III-B2).
@@
It is also noteworthy that training a separate vocoder on top of existing acoustic tokens may result in improved audio quality than the original decoded outputs, since reconstructing waveform alone may be simpler than optimizing VQ representation and reconstruction at the same time. This is exemplarily verified in AudioDec [67], MBD [94] and Vocos [95]. Therefore, some acoustic tokens directly simplify the VQ-GAN training objective back to the original VQ-VAE, where the discrete tokens are obtained first by a simple reconstruction loss, and a vocoder is trained as an additional stage, like AudioDec [67] and LSCodec [29]. These works are denoted as “VQ-VAE+GAN” in Table I.
III-A2 Latent diffusion
Different from VQ-GAN which uses GAN to generate waveforms or frequency features, some codecs also use latent diffusion [96, 97, 98] as an alternative. These codecs use discretized tokens as a condition to generate some latent acoustic space, e.g. from a pretrained continuous speech autoencoder. Since diffusion models are strong generative models, acoustic tokens of this type does not need discriminators and adversarial training like VQ-GAN. For instance, LaDiffCodec [17] uses EnCodec tokens to condition the diffusion process from Gaussian noise to the latent space in a pretrained and frozen waveform autoencoder. This is to bridge the gap of reconstruction quality between discrete and continuous representations and improve the codec performance compared to the original acoustic tokens. Inference efficiency is a major concern of these models unless specifically optimized in limited sampling steps.
@@ III-B1 Motivation
In this section, we describe the most common type of neural acoustic tokens (speech codecs) that are designed only with the objective of speech signal reconstruction. Those acoustic tokens are optimized towards better signal or perceptual quality under bitrates as low as possible.

### v3

@@ II-B3 Finite Scalar Quantization (FSQ)
The quantized values for each dimension are thus integers ranging from $\lfloor L/2\rfloor$ to $\lfloor L/2\rfloor$ [^2] Following [mentzer2024finite], this is the symmetric case for $L$ being odd. When $L$ is even, there is an offset before rounding to obtain asymmetric quantized values.. For a $d$-dimensional vector $\bm{x}$, there are $V=L^{d}$ possible quantization outcomes. STE is also applied to pass gradients. As quantization is simply done via rounding to integers, there are no explicit codebooks associated with the FSQ process.
@@ II-B4 Other VQ Tricks
Figure 3: Diagram of GVQ (left) and RVQ (right). GVQ quantizes different partitions of the input vector independently, while RVQ sequentially quantizes the residuals.
@@
    Grouped VQ (GVQ), also known as product quantization [product_quantization]. It partitions the input vector $\bm{x}$ by dimensions and applies VQ on different groups independently. Different groups can have different or shared codebooks. Each group produces a code-vector of the same dimensionality as its partition. These code-vectors are concatenated across groups to form a final output whose dimensionality equals that of $\bm{x}$.
@@
    Residual VQ (RVQ), also known as multi-stage quantization [multiple-stage-vector-quantization]. It adopts a serial approach that iteratively quantizes the residual of the last quantizer. Similar to GVQ, RVQ also has multiple quantizers. For the $i$-th quantizer $q_{i}$ with input $\bm{x}_{i}$ and output code-vector $\bm{c}_{k}$, the residual is defined as $\bm{x}_{i+1}=\bm{x}_{i}-\bm{c}_{k}$. The outputs from all $q_{i}$ are finally summed as the quantized result of $\bm{x}$. In this way, information in the codebooks is supposed to follow a coarse-to-fine order, and more details in the original $\bm{x}$ can be preserved than a plain quantizer.
GVQ and RVQ can also be flexibly combined to form GRVQ [yang2023hifi] that applies RVQ on each GVQ branch for better codebook utilization. RVQ can also be applied to FSQ [parker2024scalingtransformerslowbitratehighquality]. Note that RVQ naturally produces an order of importance in residual layers, while all quantizers in GVQ are equally important. Such order of importance can also be enforced in GVQ by a “nested dropout” trick [rippel2014learning].
II-B5 Comparisons
Compared to k-means VQ, both Gumbel VQ and FSQ avoid additional loss terms during training. However, Gumbel VQ is sensitive to the temperature parameter $\tau$ in practice [shah2024improving]. FSQ, by contrast, has a simpler design and optimization procedure, and has been reported to achieve better codebook utilization [^3] Although FSQ does not maintain an explicit codebook, utilization can still be measured over the $V=L^{d}$ possible outcomes. under large vocabulary sizes than k-means VQ [mentzer2024finite].
Nevertheless, FSQ also has certain limitations. In FSQ, as the vocabulary size $V$ follows $V=L^{d}$, $d$ is usually chosen to be small (like $d=6$ in StableCodec [parker2024scalingtransformerslowbitratehighquality]). For VQ, $V$ is not related to the code-vector dimension $d$, which can therefore be set to a wider range of values. The bottleneck of such low dimensionality might also cause FSQ to underperform in small vocabulary sizes compared to a fully-utilized VQ [mentzer2024finite]. Also, the quantization space of FSQ is strictly fixed, whereas VQ methods maintain a learnable codebook. This rigidity forces the surrounding network modules, particularly the encoder, to map the data distribution into such low-dimensional structured codes, thereby placing greater demands on model capacity.
Regarding GVQ and RVQ, the inherent ordering in RVQ provides greater flexibility than GVQ in trading off bitrate and performance. However, RVQ’s sequential nature not only prevents efficient parallelization in computation, but might also complicate optimization due to the nested STE operations.
@@ III Speech Tokenization Methods: Acoustic Tokens
Acoustic speech tokens are discrete representations derived from codec models, primarily designed for speech compression and reconstruction. The audio codec technology emerged long ago. Traditional codecs, including MP3 [rfc5219], Opus [Valin2012DefinitionOT] and EVS [dietz2015overview], typically take advantage of signal processing algorithms to improve quality and lower the bitrate.
In the deep learning era, numerous codec models based on neural networks have been developed. These models typically consist of an encoder that compresses speech signals and a decoder that reconstructs the speech signals, with a quantizer situated between the two. The quantizer is also parameterized and jointly trained with the whole network in an end-to-end manner. The codebook indices produced by the quantizer are referred to as acoustic tokens. To improve the representation ability of discrete VQ spaces and thus obtain better codec performance, RVQ, GVQ, GRVQ and FSQ tricks are commonly applied in the quantization module.
We list the VQ method, number of quantizers $Q$, frame rate $F$, vocabulary size $V$ for each quantizer, and the resulting bitrate of existing neural acoustic speech tokens in Table II in the appendix.
@@ III-A Model Architectures
[…]
Figure 5: Major generator architectures of VQ-GAN-based acoustic tokens. “Conv.”, “Q.” and “Trans.” are short for convolution, quantizer and Transformer, respectively.
@@ III-A1 VQ-GAN
[…]
@@
The CNN-based generator is the most widely used architecture so far in codec models. SoundStream [zeghidour2021soundstream] and EnCodec [encodec] are two famous early neural codec models that operate in an end-to-end VQ-GAN manner. They receive time-domain waveforms as inputs and directly reconstruct waveforms. Their encoder and decoder have a mirrored architecture to perform down and up-samplings. In SoundStream, the encoder and decoder are purely constructed by convolutional neural networks (CNNs) while EnCodec augments them with an LSTM. The CNN encoder down-samples the waveform to a high-dimensional embedding sequence, whose frame rate is determined by the sampling rate, CNN kernel sizes and strides at a fixed ratio. The continuous embeddings are passed through an RVQ quantizer. The resulting quantized vectors are summed and then passed to a CNN decoder to reconstruct the waveform. The training criteria include reconstruction loss (in the time and frequency domain), adversarial loss, feature matching loss, and quantization losses for RVQ layers. To allow for a flexible choice of bitrates, structured dropout is adopted where the number of codebooks in the RVQ module can be randomly chosen [zeghidour2021soundstream], such that only a portion of quantizers in front are activated during training. The resulting acoustic tokens can consequently reside in variable bitrates depending on the chosen number of RVQ quantizers. The inputs and outputs of the codec model can also be frequency-domain features like magnitude and phase spectra for reducing computation burden [du2024funcodec]. There, the convolution kernels are typically 2D instead of 1D in the time-domain codecs.
@@
Later, Transformers [transformer] have been adopted, e.g. in Single-Codec [singlecodec] and Mimi [kyutai2024moshi]. They can be directly applied to frequency-domain inputs and outputs. When operating on waveform-domain inputs or outputs, a CNN [kyutai2024moshi] or patchifying [wu2024ts3codectransformerbasedsimplestreaming, parker2024scalingtransformerslowbitratehighquality] operation is usually added before or after the Transformer blocks. In Mimi, a shallow Transformer layer is added after the CNN-based encoder, and vice versa in its decoder. Recently, some propose to use purely Transformer-based backbone and discard the CNN blocks, e.g. TS3-Codec [wu2024ts3codectransformerbasedsimplestreaming]. As Transformers demonstrate superior modeling ability and scaling property, these works prove to outperform CNN-based codecs either with less computation [wu2024ts3codectransformerbasedsimplestreaming] or larger scale [parker2024scalingtransformerslowbitratehighquality]. However, to ensure stream-ability, an attention mask should be employed [kyutai2024moshi]. The encoder and decoder can also be designed to be different. For example, Single-Codec [singlecodec] uses Conformer [conformer] encoder and CNN decoder, while LSCodec [guo2024lscodec] uses the reverse configuration.
In addition, U-Net-based codecs employ multiple quantizers at different layers of the network, rather than relying on a single GVQ or RVQ module for the entire codec. Typical examples in this category include CoFi-Codec [guo2024speaking] and ESC [gu2024esc]. In such designs, each sub-encoder or decoder in the U-Net can be implemented with a CNN or Transformer, offering more flexible control over the resolution of each VQ stream (Section III-B2).
@@
It is also noteworthy that training a separate vocoder on top of existing acoustic tokens may result in improved audio quality than the original decoded outputs, since reconstructing waveform alone may be simpler than optimizing VQ representation and reconstruction at the same time. This is exemplarily verified in AudioDec [audiodec], MBD [san2023discrete] and Vocos [siuzdak2024vocos]. Therefore, some codec models directly simplify the VQ-GAN training objective back to the original VQ-VAE, where the discrete acoustic tokens are obtained first by a simple reconstruction loss, and a vocoder is trained as an additional stage, like AudioDec [audiodec] and LSCodec [guo2024lscodec]. These works are denoted as “VQ-VAE+GAN” in Table II.
III-A2 Diffusion
Different from VQ-GAN which uses GAN to generate waveforms or frequency features, some codecs also use denoising diffusion [ho2020denoising, song2021scorebased] or flow matching models [lipmanflow] as an alternative. Since diffusion and flow matching belongs to the same family of generative models, we collectively refer to them as “diffusion” throughout this paper. These diffusion-based codecs use discretized tokens to condition the transformation of standard Gaussian distributions to the distribution of some continuous acoustic features, e.g. spectrogram features, or the latent space of a pretrained speech autoencoder. The diffusion loss can be propagated back to the encoder and quantizer in this design, like [liu2024semanticodec]. The encoding process in such codecs is identical to that of VQ-GAN. In the decoding process, the decoder runs an iterative sampling process to generate the target acoustic features, which are converted to waveforms by a separate pretrained model.
III-A3 Comparisons
Compared to diffusion, VQ-GAN-based codecs are more intuitive in design, and have been a well-established method. Within this category, different architectures offer distinct advantages: CNN-based models are usually lightweight and context-invariant because of a limited receptive field. Transformer-based models are easier to scale and believed to have better compression capacity. U-Net-based models offer greater flexibility in quantizer resolutions, but the correlation of tokens from different quantizers may be more complex for downstream modeling compared to adjacent quantizers in a single RVQ module.
However, VQ-GAN-based codecs rely on sophisticated discriminators, which is crucial to the performance. In contrast, diffusion-based codecs do not need adversarial training, and thus have a simpler training objective. Inference latency is a major concern of diffusion-based codec models, unless specifically optimized for limited sampling steps.
@@ III-B1 Motivation
In this section, we describe the most common type of neural acoustic tokens (speech codecs) that are designed only with the objective of speech signal reconstruction. Those acoustic tokens are designed to achieve better objective or perceptual quality at the lowest possible bitrates.